why dipole is used.can u send me the accurate definition of this.if we don't use dipole then what will happen


 

Grade:12

1 Answers

bibhash jha
14 years ago

3Consider a pair of charges $\pm e$ a small distance ${\bf s}$ apart (${\bf s}$ is the vector joining the negative charge to the positive charge and is illustrated in Figure  3.5 such that we can define

Figure 3.5: A pair of opposite charges are a distance $s$ apart. A dipole is formed by letting $s \rightarrow 0$.
\includegraphics [scale=0.7]{fundfig33a.eps}

 

\begin{displaymath} {\bf m} = e{\bf s}, \end{displaymath} (3.14)

which is called the dipole moment. The resulting dipole potential is

\begin{displaymath} 4\pi \epsilon F = -{e\over r} + {e\over \vert{\bf r} - {\bf... ...r (x^{2} + y^{2})^{1/2}} + {e\over ((x-s)^{2} + y^{2})^{1/2}}. \end{displaymath}


If we assume that $s$ tends to zero, we may expand the second term in a Taylor series for small $s$. Thus, we obtain

\begin{displaymath} 4\pi \epsilon F \approx -{e\over r} + {e\over (x^{2} + y^... ...1/2}} = {e\over r}\left [-1 + (1 + {xs\over r^{2}})\right ]. \end{displaymath}


Hence,

\begin{displaymath} 4\pi \epsilon F = {xes\over r^{3}} = {es\cos \theta\over r^{2}} = {m\cos \theta\over r^{2}}. \end{displaymath} (3.15)

The resulting electric field components are given by evaluating

${\bf E} = - \nabla F$ in spherical coordinates. Hence

$\displaystyle E_{r}$ $\textstyle =$ $\displaystyle {2m\cos \theta\over 4\pi \epsilon r^{3}}$  
$\displaystyle E_{\theta}$ $\textstyle =$ $\displaystyle {m\sin \theta \over 4\pi \epsilon r^{3}}$ (3.16)


Thus, ${\bf E}$ falls off in magnitude like $r^{-3}$ and $E_{\theta}$ vanishes when

$\theta = 0, \pi$. The electric field for a dipole is shown in Figure 3.6.

Figure 3.6: The electric field lines for an electric dipole.

 

 

 

Electric dipoles are very important . They exist everywhere around us . The presence of electric dipoles in molecules affect their chemical and physical behaviour . They also find use in antennas.

\includegraphics [scale=0.7]{fundfig34.eps}